OpenStudy (anonymous):

The vertex of the function’s graph is given. Find c. y = x2 + 10x + c; (-5, -27)

OpenStudy (mathmale):

There are several versions of the equation of a parabola (which is what you have here). One of them is the standard form y=ax^2 + bx + c. That's the form your y = x2 + 10x + c has. the other is the vertex form: \[y=x(x-h)^2+k,\]where (h,k) is the vertex (a point).

OpenStudy (mathmale):

for the same parabola, these two equations should give you one and the same graph. Thus: You could equate your y = x2 + 10x + c to the vertex form of the equation: y = x2 + 10x + c = a(x-h)^2+k. Important: Please do not write 'x2' for 'the square of x.' Instead, write x^2 or \[x ^{2}\]

OpenStudy (mathmale):

What is h? What is k? what's left to find?

OpenStudy (anonymous):

what?

OpenStudy (mathmale):

Your job is to find the value of c in \[y=x^2+10x+c.\]Where do you think that "c" value is going to come from? Why do you think I mentioned the "vertex form of the equation of a parabola?"

OpenStudy (mathmale):

Sorry. We were on the verge of answering this problem. OpenStudy tells me you're looking at some other problem now. If you're back within 2 minutes, I'd be glad to help you finish the problem at hand.

OpenStudy (anonymous):

i just guessed because i have no idea what anything you said meant. thank you for trying though

OpenStudy (mathmale):

All right. How do you pick up this info? Are in a real classroom, or are you taking an online course, or what?

OpenStudy (mathmale):

Have you seen the "vertex form of the equation of a parabola" before?

OpenStudy (anonymous):

im taking an online course and the videos that are suppose to teach us suck at explaining things

OpenStudy (mathmale):

Have you no reference book or anything like that? Nothing but videos?

OpenStudy (anonymous):

only videos.

OpenStudy (mathmale):

what a shame. I couldn't imagine learning material like this from only videos. In this problem you are told that the vertex of the graph of the given quadratic function is at (-5,-27). Would you be able to sketch a rough graph of the function (which is a parabola), based upon that information?

OpenStudy (mathmale):

By asking this question, I'm trying to assess what you know about this topic. You don't actually have to sketch the graph for me; rather, I just want to know whether your could nor not, based upon what you know.

OpenStudy (mathmale):

Sorry to learn that you're now off-line. I really wanted to help you, but found that you've apparently just disappeared without notice.